Invarianzen

Eigenschaft  top. Eigensch. erblich quot.–invar. prod.–invar.
1.Abax.  Ende Kap. I
2.Abax.  Ende Kap. I
separabel  Ende Kap. I
Zusammenhang (III.1.5) (III.1.5)
Kurvenzusammenhang (III.2.6)  (III.2.6)
Hausdorffsch  (III.3.7)  (III.3.6)  (III.3.8)
Regularität  (III.4.4)  (III.4.5)  (III.4.6)
Normalität  (III.5.5)
Quasikompaktheit  (V.1.8)  (V.1.9)  (V.1.10)